OpenStudy (anonymous):

Simplify: ( 5+ 2√7)^2 is the answer 25+ 4√7?

Parth (parthkohli):

Use this formula/identity. \(\Large \color{Black}{\Rightarrow (a + b)^2 = a^2 + 2ab + b^2}\)

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow 5^2 + (2*5*2\sqrt{7}) + (2\sqrt7)^2 }\)

OpenStudy (anonymous):

I'm sorry it's just that I'm really bad at multipying or adding radicals :/

Parth (parthkohli):

Ok :D \(\Large \color{Black}{\Rightarrow 25 + 20\sqrt7 + 28 }\) Wait a min...I'll teach you how I got this.

Parth (parthkohli):

Ok do you wanna twid?

OpenStudy (anonymous):

sure :)